OpenStudy (anonymous):

write quadratic equation -10 and -4

OpenStudy (unklerhaukus):

\[x^2+14x+40=(x+10)(x+4)\]

OpenStudy (saifoo.khan):

x = -4 x = -10 x +4 =0 x + 10 = 0 (x+4)(x+10) = 0

OpenStudy (anonymous):

Put it back into it's unsolved form. x = -10 x + 10 = 0 x = -4 x + 4 = 0 (x + 10)(x + 4) Can you do it now?